OpenStudy (anonymous):

compute the discriminant of 2x^2-4x+9=2

OpenStudy (luigi0210):

If I'm right the discriminant is: \[\sqrt{b^2-4(ac)}\]

OpenStudy (anonymous):

oh yeah that's it

OpenStudy (anonymous):

do you know then how to describe the solutions

OpenStudy (luigi0210):

So just plug in values:\[\sqrt{(-4)^2-4(2)(7)}\]

OpenStudy (luigi0210):

and you have to set the original equal to 0: \[2x^2-4x+7=0\]

OpenStudy (anonymous):

how do I use \[\sqrt{-40}\] to describe a solutions

OpenStudy (luigi0210):

when it's negative that means that there are no REAL solutions.. there are gonna be imaginary solutions

OpenStudy (anonymous):

ok thanksss

OpenStudy (luigi0210):

Np :P